How long will it take $46,000 placed in a savings account at 10% interstate's to grow into $71,000?

a. 3.55 years

b. 6.55 years

c. 5.55 years

d. 4.55 years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

71,000=46,000*(1.1)^n

(71,000/46,000)=(1.1)^n

Taking log on both sides;

log (71,000/46,000)=n*log 1.1

n=log (71,000/46,000)/log 1.1

=4.55 years(Approx).